### v1

# Learning Fricke signs from Maass Form Coefficients

Abstract.
In this paper, we conduct a data-scientific investigation of Maass forms. We find that averaging the Fourier coefficients of Maass forms with the same Fricke sign reveals patterns analogous to the recently discovered “murmuration” phenomenon, and that these patterns become more pronounced when parity is incorporated as an additional feature. Approximately $43\%$ of the forms in our dataset have an unknown Fricke sign. For the remaining forms, we employ Linear Discriminant Analysis (LDA) to machine learn their Fricke sign, achieving $96\%$ (resp. $94\%$) accuracy for forms with even (resp. odd) parity. We apply the trained LDA model to forms with unknown Fricke signs to make predictions. The average values based on the predicted Fricke signs are computed and compared to those for forms with known signs to verify the reasonableness of the predictions. Additionally, a subset of these predictions is evaluated against heuristic guesses provided by Hejhal’s algorithm, showing a match approximately $95\%$ of the time. We also use neural networks to obtain results comparable to those from the LDA model.
KHL was partially supported by a grant from the Simons Foundation (#712100).
DLD was supported by the Simons Collaboration in Arithmetic Geometry, Number Theory, and Computation via the Simons Foundation grant 546235.
YDQ was supported by the NSF grant PHY-2019786 (the NSF AI Institute for Artificial Intelligence and Fundamental Interactions).
1. Introduction
The purpose of this paper is to study Maass forms using various techniques from machine learning. This builds on previous work for $L$-functions attached to number fields and arithmetic curves [8, 9, 10]. Attempts to interpret this prior work led to the discovery of so-called murmurations, which are statistical correlations between the root numbers of $L$-functions and their Dirichlet coefficients, first observed in the context of elliptic curves [11]. Murmurations for Maass forms were established in [1].
We show that supervised machine learning techniques can be used to predict the Fricke sign of a Maass form based on its coefficients, without indicating the level. In contrast to the papers cited previously, we note that unsupervised techniques such as $k$-means clustering and PCA were unsuccessful at producing clusters separated by Fricke sign. The relationship between murmurations of Maass forms and machine learning their Fricke sign is analogous to the relationship between murmurations of elliptic curves and machine learning their ranks (as in [10, 11]).
The machine learning experiments presented here use the database of Maass forms from the LMFDB [15]. Among the 35,416 Maass forms in the LMFDB, some 15,423 of them lack rigorously computed Fricke signs, though it is possible to guess them heuristically. The computations of Maass forms underpinning the LMFDB rely on a combination of different techniques, including automorphic certification [3], explicit trace formulas [16], and an implementation of Hejhal’s algorithm [12] described in forthcoming work [13]. After that, the Fricke sign is subsequently determined by the signs of the coefficients $a_{p}$ for each prime $p$ dividing the level of the Maass form. In practice, however, it is often quite hard to compute the Maass form eigenvalue and coefficients with sufficient precision to directly recover the Fricke sign. This difficulty explains why a substantial portion of the dataset currently has unknown Fricke signs.
Consequently, it is surprising that our machine learning results achieve high accuracy ($>96\%$) in predicting the Fricke sign. Our approach fundamentally differs from conventional methods: we consider the dataset collectively and employ machine learning tools instead of relying solely on the intrinsic properties of individual forms. Moreover, we apply the trained machine learning model to Maass forms with unknown Fricke signs in order to predict them. By comparing these predictions with heuristic guesses, we demonstrate that our results are reasonable and may provide valuable information that could help determine the signs precisely.
In future work, it will be beneficial to interpret the machine learning results, to identify which features contribute most significantly to the high accuracy, and to understand how they do so. This analysis may lead to a deeper understanding of the Fricke sign. Additionally, we anticipate that other important quantities in number theory can also be effectively approached through machine learning. We hope this methodology will open new avenues for studying various objects that are difficult to compute in conventional ways.
We conclude this introduction with a summary of each section. In Section 2.1, we review the necessary definitions. In Section 2.2, we observe murmurations of Maass forms by averaging their coefficients. In Section 2.3, we use linear discriminant analysis (LDA) to predict the Fricke sign based upon these same coefficients. In Section 2.4, we outline a strategy for predicting the Fricke sign of a Maass form based upon factorising its level, and confirm that this approach is not what is taking place in the LDA. In Section 2.5, we train a neural network on the coefficients and the spectral parameter and achieve an accuracy comparable to that attained with LDA. In Section 2.6, we compare our predictions to those based on a heuristic version of Hejhal’s algorithm.
2. Experiments with Maass forms
2.1. Definitions
For a broad overview, the reader is referred to [5]; a more extensive (but less approachable) reference is [4]. Let $\Delta$ denote the Laplace–Beltrami operator on the upper half-plane. A weight 00 Maass cuspform $f$ on $\Gamma_{0}(N)$ is a smooth square-integrable function $f:\mathcal{H}\rightarrow\mathbb{C}$ satisfying $f(\gamma z)=f(z)$, for all $\gamma\in\Gamma_{0}(N)$, and $(\Delta-\lambda)f(z)=0$, for some $\lambda\in\mathbb{C}$. We refer to $N$ as the level of $f$. Assuming the Selberg eigenvalue conjecture, we may write $\lambda=\frac{1}{4}+R^{2}$ for $R\in\mathbb{R}$. In what follows, we refer to $R$ as the spectral parameter.
If $f$ is a Maass form on $\Gamma_{0}(M)$ and $M$ divides $N$, then, for any $k$ dividing $N/M$, $f(kz)$ is a so-called oldform on $\Gamma_{0}(N)$. We will consider only Maass forms that are not oldforms, known as newforms, which live naturally on $\Gamma_{0}(N)$. We write $S_{0}^{\textup{new}}(\Gamma_{0}(N))$ for the set of weight 00 Maass newforms on $\Gamma_{0}(N)$.
It is known that $f\in S_{0}^{\textup{new}}(\Gamma_{0}(N))$ has a Fourier expansion of the form
(2.1) $f(x+iy)=\sum_{n\neq 0}a_{n}\sqrt{y}K_{iR}(2\pi|ny|)\exp(2\pi inx),$
[…]
$f(z)=w_{N}f\left(-\frac{1}{Nz}\right),$
for $w_{N}\in\{\pm 1\}$. We refer to $w_{N}$ as the Fricke sign.
There is an involution on $\mathcal{H}$ given by reflection in the imaginary axis, $z\mapsto-\bar{z}$. We call a Maass form even (resp. odd) if $f(-\bar{z})=f(z)$ (resp. $-f(z)$). Let $\sigma(f)$ be 00 (resp. $1$) if $f$ even (resp. odd). Applying equation (2.1) for a Maass form $f$ with fixed parity, we deduce:
(2.2) $f(x+iy)=\sum_{n=1}^{\infty}a_{n}\sqrt{y}K_{iR}(2\pi\lvert ny\rvert)\begin{cases}2\cos(2\pi inx),&\sigma(f)=0,\\ 2i\sin(2\pi inx),&\sigma(f)=1.\end{cases}$
The $L$-function associated to $f$ is given by $L_{f}(s)=\sum_{n=1}^{\infty}a_{n}n^{-s}$, which converges for $\mathrm{Re}(s)>1$. The completed $L$-function is:
$\Lambda_{f}(s)=\left(\frac{\sqrt{N}}{\pi}\right)^{s}\Gamma\left(\frac{s+\sigma(f)+iR}{2}\right)\Gamma\left(\frac{s+\sigma(f)-iR}{2}\right)L_{f}(s).$
The completed $L$-function satisfies the functional equation
$\Lambda_{f}(s)=\epsilon\bar{\Lambda}_{f}(1-s),$
where $\epsilon$ is the root number. In particular, we have $\epsilon=(-1)^{\sigma(f)}w_{N}$ (see [4, §8]).
Given complete information about the coefficients, the Fricke sign is easily computable; on $\Gamma_{0}(N)$ with $N$ squarefree, the coefficient $a_{N}$ encodes the Fricke sign. However, explicitly computing Maass form coefficients is extremely hard in practice. Conjecturally, all data associated to a generic Maass form is transcendental and independent of typical transcendental constants (cf. [2]).
The Fricke signs in the LMFDB were rigorously computed from the coefficients, and are missing when these coefficients were not computed with sufficient precision. The challenge most often starts with numerical difficulties in the explicit trace formula [16]. Increasing the precision of the trace formula requires computing many class numbers rigorously, which is very hard. A different approach would be to implement the confirmation algorithm from [3] for general level. Whilst there are other methods to determine the Fricke sign of a Maass form, none have been implemented rigorously.
In the following subsections, we will show that one can use methodology from machine learning to predict the Fricke sign from finitely many coefficients of a Maass newform. This is analogous to machine learning the rank (parity) of an elliptic curve, which is also manifest in the sign of a completed $L$-function [10, 11]. Furthermore, this is connected to murmurations of Maass forms as in [1], which amount to subtle statistical correlations between the sign and the coefficients. Finally, we will compare our predictions with the heuristic results from Hejhal’s algorithm described just above.
2.2. Averaging
Our analysis uses a dataset $\mathcal{L}$ containing the 35,416 rigorously computed Maass forms from the LMFDB [15]. The dataset contains the first $1,000$ Fourier coefficients $a_{n}$ for every Maass form, each of which has weight 00, trivial character, and integral level $N$ in the range from $1$ to $105$. In the analysis that follows, we are particularly interested in the parity $\sigma$ and the Fricke sign $w$ and (we will often drop the subscript from $w_{N}$ if there is no need to specify $N$). As discussed above, in the LMFDB, the value of the Fricke sign is not always rigorously computed, in which case it is given the value $w=0$. The breakdown of the Maass forms dataset $\mathcal{L}$ with different values for $\sigma$ and $w$ are given in Table 2.1.
$w=-1$ $w=1$ $w=0$ (unknown) Total
$\sigma(f)=0$ (even) $5009$ $7171$ $6173$ $18353$
$\sigma(f)=1$ (odd) $3724$ $4089$ $9250$ $17063$
Total $8733$ $11260$ $15423$ $35416$
Table 2.1. Counts of Maass forms by parity and Fricke sign.
We may write $\mathcal{L}$ as a disjoint union $\mathcal{L}_{0}\coprod\mathcal{L}_{1}\coprod\mathcal{L}_{-1}$, in which $\mathcal{L}_{i}=\{f\in\mathcal{L}:w=i\}$. Figures 2.1 and 2.2 provide clear evidence of murmuration-like distinction between the forms in $\mathcal{L}_{1}$ and those in $\mathcal{L}_{-1}$. More precisely, in Figure 2.1, for primes $p<1000$, we plot the average value of $a_{p}$ over $\mathcal{L}_{1}$ and $\mathcal{L}_{-1}$, where $a_{p}$ is as in equation (2.2). We note that the separation is much better when also taking symmetry into account. This is equivalent to the idea of separating by root number. Because the root number $\epsilon=(-1)^{\sigma(f)}w$ we explore normalizing the coefficients by multiplying by $(-1)^{\sigma(f)}$. In Figure 2.2 we see this produces an even more distinctive murmuration pattern.
Figure 2.1. Average value of $a_{p}$ over Maass Forms with given Fricke sign, with and without separating by symmetry
Figure 2.2. Average value of $(-1)^{\sigma(f)}a_{p}$ over Maass Forms separated by Fricke sign.
2.3. Predicting the Fricke sign with LDA
The clear separation in Fricke sign provided by averaging $(-1)^{\sigma(f)}a_{p}$ indicates that it may be possible to predict the Fricke sign based on these features, perhaps even using a fairly simple technique. We choose to undertake a Linear Discriminant Analysis (LDA) to learn a linear decision boundary between classes of Fricke sign. LDA is a supervised dimensionality reduction and classification technique based on Bayes theorem [7, Section 4.3].
By writing each $f\in\mathcal{L}$ as in equation (2.2), we may construct the following $1000$-dimensional feature vector:
$\mathcal{D}=\left\{\left((-1)^{\sigma(f)}a_{n}\right)_{i=1}^{1000}:f\in\mathcal{L}\right\}\subset\mathbb{R}^{1000}.$
We may write $\mathcal{D}=\mathcal{D}_{0}\coprod\mathcal{D}_{1}\coprod\mathcal{D}_{-1}$, where $\mathcal{D}_{i}$ contains only vectors corresponding to forms in $\mathcal{L}_{i}$.
We ran supervised learning experiments on the labeled dataset $\mathcal{D}_{1}\coprod\mathcal{D}_{-1}$. In each experiment, we split a subset into testing, training, and validation sets, and the testing set accounted for $20\%$ of the relevant data with the training-validation split also 80-20. Firstly, we trained the LDA using all available training data ($12795$ observations) and recorded $96.1\%$ accuracy on the validation data. Secondly, we masked the training data so as to include only even forms ($7772$ observations) and recorded $94.9\%$ accuracy on the similarly masked validation data. Thirdly, we trained only data from odd forms ($5023$ observations), and recorded $96.3\%$ accuracy on the similarly masked validation data. In all three cases, the accuracy on the testing set is almost identical to that on the validation set. These experiments indicate that, without any hyperparameter tuning, we are able to get high predictive accuracy.
Building on the paragraph above, we also make predictions using the trained LDA model for the Maass forms with unknown Fricke sign, that is, those in $\mathcal{L}_{0}$. Then, in order to check whether our predictions are reasonable, we examined if the average value of $(-1)^{\sigma(f)}a_{p}$ for the newly predicted Fricke sign follows similar murmuration patterns as the rigorously proved (known) values. Figures 2.3 and 2.4 demonstrate that our predicted values mimic the murmuration patterns seen in Figure 2.1. By comparing Figures 2.3 and 2.4 we make the interesting observation that, when $p$ is small, there is greater similarity between genuine and predicted murmurations for odd Maass forms than there is for even Maass forms.
Figure 2.3. Average value of $a_{p}(-1)^{\sigma(f)}$ for Maass Forms with odd parity, separated by Fricke sign.
Figure 2.4. Average value of $a_{p}(-1)^{\sigma(f)}$ for Maass Forms with even parity, separated by Fricke sign.
2.4. Embedding the Fricke signs into the Fourier coefficients
The (global) Fricke sign introduced in Section 2.1 may be written as a product of the form
(2.3) $w_{N}=\prod_{p|N}w_{p},$
in which, for a prime number $p$, the number $w_{p}$ is the local Fricke sign. If $p$ divides the level $N$ then we have
(2.4) $a_{p}=\frac{-w_{p}}{\sqrt{p}}$
We conclude from equation (2.4) that the Fricke sign is determined by the sequence $(\mathrm{sgn}(a_{p}))$ of signs and the number of prime factors of $N$.
If the Fricke sign is unknown, we do not know the value of $a_{p}$ for $p|N$, and these values are defined to be 00 in the LMFDB. Similarly, when $\gcd(n,N)>1$, we have $a_{n}=0$. Given that we are training on a dataset where the Fricke sign is known, one might wonder whether perhaps LDA is just learning how to determine the Fricke sign from the $a_{n}$ with $\gcd(n,N)>1$. To clarify, we do experiments that show this is not the case. Before describing the experiments, we note that, for level $1$ Maass forms (of which there are $2202$ in our dataset), as $\gcd(n,1)=1$, the Fourier coefficients do not directly encode information about the Fricke sign in the above way. Similarly, as the largest level in our dataset is $105$, for any of the primes $p>105$, the value of $a_{p}$ does not directly encode this information.
In our previous section we used vectors of the form $(a_{n})_{n=1}^{1000}$ to make predictions with LDA. In this section, we will explore, firstly, the impact of only using coefficients with prime index, and, secondly, the impact of removing the extra information about the Fricke sign embedded in the coefficients when $\gcd(n,N)>1$ by setting all of these coefficients to zero. Subsequently, we examine the accuracy of both the first and second experiments based on the number of coefficients used to train the data.
In Figure 2.5, we compare the average value of $(-1)^{\sigma(f)}a_{p}$ and $(-1)^{\sigma(f)}a^{\prime}_{p}$ over $f\in\mathcal{L}_{1}$ and $f\in\mathcal{L}_{-1}$, where
$a^{\prime}_{n}=\begin{cases}0,&\gcd(n,N)>1,\\ a_{n},&\gcd(n,N)=1.\end{cases}$
This matches the data that we are trying to predict, as all $a_{n}=0$ if $\gcd(n,N)\neq 1$ for the Maass forms with unknown Fricke sign. For $p>105$ we have $a_{p}=a^{\prime}_{p}$, and so we restrict the graph to the relevant primes. We see that there is very little impact to the average, and still good separation between the different Fricke signs.
Figure 2.5. Comparing average values of $a_{p}$ versus $a^{\prime}_{p}$ by Fricke sign
In Table 2.2, we record the accuracy of LDA when applied to vectors of the following form:
$(a_{n})_{n=1}^{1000},~{}(a^{\prime}_{n})_{n=1}^{1000},~{}(a_{p_{i}})_{i=1}^{168},~{}(a^{\prime}_{p_{i}})_{i=1}^{168},$
where $p_{i}$ is the $i$th prime (there are $168$ primes less than $1000$). In these experiments, we see that replacing certain coefficients by zero does not make much difference to the accuracy. This is good news; we believe this means LDA should do well at predicting the Fricke sign in the case where it is unknown. Since the values of the coefficients are multiplicative, we would expect the $a_{p}$ to suffice. It is therefore surprising that there is a substantial improvement in our accuracy when using all $a_{n}$ rather than just $a_{p}$. We do not yet have a complete understanding of why this is true. It could be that the $a_{n}$ see the smaller primes more often due to their multiplicative nature, and these have a stronger impact on accuracy.
Features all parity (normalized) even parity odd parity
$a_{n}$ 0.9612 0.9488 0.9633
$a^{\prime}_{n}$ 0.9456 0.9564 0.9633
$a_{p_{i}}$ 0.8625 0.8261 0.8800
$a^{\prime}_{p_{i}}$ 0.8615 0.8329 0.8769
Table 2.2. Comparing the accuracy of LDA predictions for different features and parities.
We next analyze how these methods compare with each other when predicting the Fricke sign of Maass forms in $\mathcal{L}_{0}$. Predictions based on $(a_{n})_{n=1}^{1000}$ and $(a_{p_{i}})_{i=1}^{168}$ agree $83.29\%$ of the time. Predictions based on $(a_{n})_{n=1}^{1000}$ and $(a^{\prime}_{n})_{n=1}^{1000}$ agree $97.50\%$ of the time. Predictions based on $(a_{p_{i}})_{i=1}^{168}$ and $(a^{\prime}_{p_{i}})_{i=1}^{168}$ agree $99.81\%$ of the time. This is consistent with the accuracy results outlined above in that zeroing out the coefficients $a_{n}$ when $\gcd(n,N)>1$ does not have much impact on the predictions, but using $a_{n}$ versus $a_{p}$, does have a substantial impact on the predictions.
Finally, we analyze the different LDA methods used in this section and determine their accuracy relative to the number of coefficients used to train the data; see Figure 2.6. We again notice that LDA using $(a_{n})_{n=1}^{1000}$ has better accuracy than just using $(a_{p_{i}})_{i=1}^{168}$, and that there is little difference across the board for correct coefficients versus zeroed out coefficients. Interestingly, we note that there is striking improvement in accuracy for including more $a_{n}$ with small $n$, but then a leveling off as $n$ increases. This still occurs but is less pronounced when using $a_{p}$.
Figure 2.6. Accuracy on validation set as number of $a_{n}$ increases
2.5. Predicting the Fricke sign with a neural network
Motivated by the success of LDA, we trained neural networks to predict the Fricke sign using feature vectors of the form $(a_{2},a_{3},\dots,a_{p_{d}},R)$. As in the previous experiments, we take $d=168$, so as to use all primes $<1000$. In fact, we construct two different neural networks, one for each possible parity of the form. This is motivated by Figure. 2.1. Unlike in the previous sections, we have included the spectral parameter $R$ as a feature. Since the analytic conductor of a Maass form is (roughly) equal to $\frac{NR^{2}}{4\pi^{2}}$, experiments using $R$ as a feature alongside the coefficients are similar to the experiments in [6] (in which the elliptic curve conductor is used alongside the Frobenius traces to predict the rank). We note that neglecting $R$ significantly reduced the accuracy of the predictions (cf. Figure 2.10).
The neural network architecture is shown in Figure 2.7. We train the networks over $4\times 10^{4}$ iterations, using Adam optimization with learning rate $10^{-3}$. This achieves around $95\%$ accuracy for even forms and $94\%$ accuracy for odd forms. For the loss function, we use binary cross-entropy, with the classes corresponding to two different possibilities for the Fricke sign $w_{N}$.
$\tiny a_{2}$ $a_{3}$.. $a_{p_{d}}$ $R$$\tiny\mathsf{n}_{1}^{(1)}$ $\tiny\mathsf{n}_{2}^{(1)}$ $\tiny\mathsf{n}_{3}^{(1)}$ $\tiny\mathsf{n}_{4}^{(1)}$ $\tiny\mathsf{n}_{5}^{(1)}$ $\mathsf{n}_{1}^{(2)}$ $\mathrm{Prob}(w_{N}=1)$$\mathrm{ReLu}$$\sigma$
Figure 2.7. Neural network architecture used for predicting the Fricke sign from feature vectors of the form $(a_{2},a_{3},\dots,a_{p_{d}},R)$ where $p_{i}$ denotes $i^{\mathrm{th}}$ prime number, and $R$ denotes the spectral parameter. Here $\mathsf{n}_{i}^{(j)}$ denotes $i^{\mathrm{th}}$ node in $j^{\mathrm{th}}$ layer. The spectral parameter $R$ is switched off for generating one of the visualizations in Figure 2.10.
The saliency analysis of the trained neural network (shown in Figure 2.8) indicates the coefficients $\{a_{p}~{}|~{}p\leq 53\}$ and the spectral parameter $R$ contribute the most to the decision of the neural network. While the input parameters, including the coefficients $(a_{p_{i}})_{i=1}^{d}$ and the spectral parameter $R$, are normalized to have unit variance over the dataset, the standard deviations of the distributions of the coefficients indicate a sharp jump around $p=53$, as shown in Figure 2.9. This variance is caused by the $a_{p}$ having only three distinct values in our dataset, $\{0,\pm 1/\sqrt{p}\}$, when $p$ divides the level. This indicates that it is possible that the neural network architecture is making more use of the way the Fricke sign is embedded in the the Fourier coefficients when $p|N$. The Maass forms in the LMFDB have analytic conductor up to about $600$. Motivated by this observation, we measure the influence that the number of the coefficients has on the performance of the network. In particular, we train each neural network on a subset of coefficients of varying size, and calculate the resulting accuracy under the same choice of hyperparameters. The dependence of the accuracy on the number of coefficients is shown on Figure 2.10.
Figure 2.8. Saliency map of a neural network trained for predicting Fricke sign.
Figure 2.9. Variances of the coefficients $a_{p}$ for Maass forms from [15].
Figure 2.10. Dependence of the accuracy on the number of the coefficients $a_{p}$ (for $p$ prime) used as the input to the neural network trained for predicting Fricke sign. Here $R$ denotes the spectral parameter.
2.6. Comparison of Hejhal’s algorithm, LDA, and neural networks
The rigorous computations of Maass forms in the LMFDB were computationally expensive. It is easier to make heuristic guesses at the Fricke signs. One approach is to modify (the original, heuristic, version of) Hejhal’s algorithm [12] to apply to general level. This works by first guessing an eigenvalue and a set of Atkin–Lehner involution signs. Then, one evaluates a Maass form $f$ with that eigenvalue and Atkin–Lehner behavior at several points, and also evaluates $f$ at images of these points under $f(z)\mapsto f(\gamma z)$ for $\gamma\in\Gamma_{0}(N)$. If the guessed eigenvalue is close enough to be correct, and if the Atkin–Lehner signs are correct, then these expansions must approximately agree by the automorphy of $f$. By truncating the Fourier expansions, this becomes an over-determined, approximate, homogenous linear system in the coefficients. The non-linearity of the group action non-trivially mixes the coefficients to make the system of equations solvable in practice.
We ran this heuristic Hejhal’s algorithm as given in [14] for each of the Maass forms in the LMFDB with unknown Fricke sign. As the Fricke signs were unknown, it was necessary to try every possible combination of Atkin–Lehner signs, which makes the computation much more expensive. For each eigenvalue and set of signs, Hejhal’s algorithm produces a candidate list of coefficients. We can heuristically guarantee that these coefficients are close to correct if they satisfy Hecke relations and are close to known coefficients.
This heuristic algorithm yields expansions and Fricke signs that are heuristically extremely close to correct for 4,595 of the 15,423 Maass forms in the LMFDB with unknown Fricke signs. We consider these as “probably correct” and compare the models here against this data.
Remark 2.1.
Hejhal’s intention was to iterate the algorithm to find spectral eigenvalues and their Maass forms. For each choice of Atkin–Lehner signs, we can regard Hejhal’s algorithm as a demanding root-finding algorithm in the unknown spectral eigenvalue. For those forms with unknown Fricke sign, it would be necessary to repeatedly run this algorithm with every combination of Atkin–Lehner signs and with several eigenvalue candidates. It was surprising to the authors that the current precision in the LMFDB isn’t sufficient for even heuristic versions of Hejhal’s algorithm to converge in practice.
In Table 2.3, we compare the predictions using the 4 different LDA methods from Table 2.2 with those from the heuristic Hejhal algorithm. The features that achieve the greatest agreement are the standard Fourier coefficients $a_{n}$. The performance on odd Maass forms is slightly better than on even ones, similar to results on the validation set. The high accuracy is sustained when we restrict to forms of fixed parity. For example, using $a_{n}$ as features, we achieve $95.45\%$ accuracy on all forms, and, if one only looks at odd (resp. even) forms, then the accuracy is $96.09\%$ (resp. $94.61\%$). A comparison between LDA and Hejhal’s heuristic for Maass forms with fixed parity is documented in Table 2.4. In Table 2.5, we compare the predictions using neural networks with those from the heuristic Hejhal’s algorithm.
[…]
$a_{n}$ 95.45
$a^{\prime}_{n}$ 94.84
$a_{p_{i}}$ 82.93
$a^{\prime}_{p_{i}}$ 82.87
Table 2.3. Comparing the LDA predictions with the heuristic approach based on Hejhal’s algorithm.
[…]
odd 96.09 (1877 out of 1984)
even 94.61 (2509 out of 2611)
Table 2.4. Comparison of LDA with the heuristic approach based on Hejhal’s algorithm for Maaass forms with fixed parity. For LDA, we used feature vectors of the form $(a_{n})_{n=1}^{1000}$.
[…]
both $87.37$
odd $85.79$
even $89.46$
Table 2.5. Comparing the neural network predictions with the heuristic approach based on Hejhal’s algorithm.